OpenStudy (help_me_pls):

Which expression is equal to f(x) + g(x) given f(x) = x² + 10 and g(x) = x – 3 ? a. 2x² + 7 b. 7x² + x - 3 c. x² + x + 7 d. x² - x - 10

OpenStudy (calculusxy):

You would just to combine the two functions of \(f(x)\) and \(g(x)\). \(f(x) = x^2 + 10\) \(g(x) = x - 3\) \(\bf \color{red}{f(x)} + \color{green}{g(x)} = \color{red}{x^2 + 10} + \color{green}{x-3}\) Now you just need to combine the like terms to simplify the expression.

OpenStudy (help_me_pls):

I think that it's A but I also think it may be C

OpenStudy (mathmale):

Please share your thoughts. What makes you think (tentatively) that "it's A" and "it may be C?" Show your work, please. Then I could give you meaningful feedback.

OpenStudy (calculusxy):

It cannot be A because there's no two \(x^2\) terms that you can combine to make \(2x^2\). Here's how I used to combine like terms especially when there were different types of variables raised to different powers.